OpenStudy (anonymous):

When a capacitor has a charge of absolute value 80 µC on each plate the potential difference across the plates is 10 V. How much energy is stored in this capacitor when the potential difference across its plates is 120 V?

OpenStudy (anonymous):

Do you know how to find the capacitance when charge and voltage are given?

OpenStudy (anonymous):

Anyway, use q=C*V to find C. Plug it into E=1/2*C*V^2.

OpenStudy (anonymous):

are you sure from this laws :D I applied this rules and the answer is wrong :S

OpenStudy (anonymous):

0.48 mJ 58 mJ 0.12 J 0.90 GJ 0.36 J

OpenStudy (anonymous):

@pully, are these the options?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when i followed @arcticf0x's method, i got 57.6 mJ

OpenStudy (anonymous):

I got 5.76x10^10

OpenStudy (anonymous):

so it will be 58 mj

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

i don't understand the answer i have the same question

OpenStudy (anonymous):

Apply this Ahmed q=C*V to find C. Plug it into E=1/2*C*V^2. Then you will get 57.6 mJ